Grade 12th passIIT JEE Entrance Exam

A ball is thrown upward at t = 0 sec, it is at the same height at t = 1 sec and t = 3 sec then maximum height attained by the ball is
1)10m
2)20m
3)30m
4)40m

Anup singh
8 Years agoGrade 12th pass

1 Answer

Sarfaraz Ahamed
8 Years ago
The Answer is 2)20m
let us consider that the ball is at a height h at time t=1s and t=3s
let us assume that g=10m/s2
then h=u*1-1/2(g*12) for t=1s
and h=u*3-1/2(g*32) for t=3s
therefore h=u-5 for t=1s and h=3u-45 for t=3s
therefore u-5=3u-45
therefore u=20m/s
the maximum height is given by u2/(2g)=202/2*10=20m